OpenStudy (btaylor):

Rationalizing the denominator: is it possible to do this one in 1 step? \[1 \over 1 + \sqrt{3} - \sqrt {5}\]

OpenStudy (btaylor):

I did this: \[\frac{1}{1 + \sqrt{3} - \sqrt{5}} \times \frac{-1+\sqrt{3}-\sqrt{5}}{-1 + \sqrt{3}- \sqrt{5}} = \frac{-1+\sqrt{3}-\sqrt{5}}{7-2\sqrt{15}}\] Is there a simpler way?

OpenStudy (btaylor):

@jackellyn @ajprincess help please!

OpenStudy (jackellyn):

It looks fine to me, i'm not sure any other way. Sorry.

OpenStudy (ajprincess):

ya I too am nt sure of any other way..sorry.

OpenStudy (anonymous):

let \[a=\sqrt{3}-\sqrt{5}\] \[1/(1+a)\] \[1/(1+a)*(1-a)/(1-a)\] \[(1-a)/(1-a ^{2})\]

OpenStudy (anonymous):

or \[1/(1+\sqrt{3}-\sqrt{5})*(1+\sqrt{3}+\sqrt{5})/(1+\sqrt{3}+\sqrt{5})\]

OpenStudy (btaylor):

thanks!